Primus, Inc., owns all outstanding stock of Sonston, Inc. For the current year, Primus reports net income (exclusive of any investment income) of $488,000. Primus has 50,000 shares of common stock outstanding. Sonston reports net income of $88,000 for the period with 50,000 shares of common stock outstanding. Sonston also has 10,000 stock warrants outstanding that allow the holder to acquire shares at $13.50 per share. The value of this stock was $27 per share throughout the year. Primus owns 4,000 of these warrants.

What amount should Primus report for diluted earnings per share? (Round your intermediate percentage value to the nearest whole number and the final answer to 2 decimal places.)
Answer #1

Sonston's net income

88000

Shares:

Outstanding

50000

Assumed Conversion of Stock Warrants

10000

Repurchase of Treasury Stock with Proceeds of Stock Warrants (10,000 x $13.50 = $135,000/$27)

(5000

5000

Shares For Primary Earnings Per Share Computation

55000

Shares Controlled by Primus: 50,000 + (40% of 5,000) = 52000

Percentage of Total Held by Primus: 52000/55000 = 95% (rounded)

     

Income to be Included in Consolidated Diluted Earnings Per Share = $88,000 x 95% = $83200

Total income = Net Income – Primus+ Net Income included from Sonston = 488000+83200 = $571200

Outstanding Shares of Primus =50000

CONSOLIDATED DILUTED EARNINGS PER SHARE = $571200/50,000 = $11.42